OpenStudy (anonymous):

Please help I don't understand the table shows how the number of sit-ups marla does each day has changed over time. at this rate how many sit-ups will she do on day 12? Explain your steps in solving this problem day 1 day 2 day 3 day 4 day 5 17 21 25 29 33

OpenStudy (texaschic101):

what you have is an arithmetic series. There is a formula for this... an = a1 + (n - 1) d a1 = first term, so it would be 17 d = the common difference, so it would be 4 n = the number of the term you want to find, so it would be 12. now we sub... a12 = 17 + (12 - 1) 4 a12 = 17 + (11)(4) a12 = 17 + 44 a12 = 61 your 12th term is 61 read more about this type of problem here... http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

no problem :)